y^2+2x-4y+9=0 find the standard form
Answer by richwmiller(17219)

This is a parabola that opens on the ;left because the x and y are switched.
2x+y^2+9 = 4y
2x=-y^2+4y-9
x=-y^2/2+2y-9/2
not sure which you want.